### v2

# Target Pebbling in Trees

Abstract
Graph pebbling is a game played on graphs with pebbles on their vertices. A pebbling move removes two pebbles from one vertex and places one pebble on an adjacent vertex. A configuration $C$ is a supply of pebbles at various vertices of a graph $G$, and a distribution $D$ is a demand of pebbles at various vertices of $G$. The $D$-pebbling number, $\pi(G,D)$, of a graph $G$ is defined to be the minimum number $m$ such that every configuration of $m$ pebbles can satisfy the demand $D$ via pebbling moves. The special case in which $t$ pebbles are demanded on vertex $v$ is denoted $D=v^{t}$, and the $t$-fold pebbling number, $\pi_{t}(G)$, equals $\max_{v\in G}\pi(G,v^{t})$. It was conjectured by Alcón, Gutierrez, and Hurlbert that the pebbling numbers of chordal graphs forbidding the pyramid graph can be calculated in polynomial time. Trees, of course, are the most prominent of such graphs. In 1989, Chung determined ${\pi}_{t}(T)$ for all trees $T$. In this paper, we provide a polynomial-time algorithm to compute the pebbling numbers ${\pi}(T,D)$ for all distributions $D$ on any tree $T$, and characterize maximum-size configurations that do not satisfy $D$.
Keywords: graph pebbling, target pebbling, paths, trees
MSC 2020: 05C57 (Primary), 05C05, 90B10, 91A43 (Secondary)
1 Introduction
Graph pebbling is a mathematical game or puzzle that involves moving pebbles along the edges of a connected graph, subject to certain rules. The objective of the game is to place a certain number of pebbles on specific vertices of the graph from any sufficiently large configuration of pebbles. The original application of graph pebbling involved only the case of placing a single pebble on a certain target. As a method to prove results in this realm, the problem was necessarily generalized in [6] to place multiple pebbles on a single target. Similarly, the problem was generalized in [7, 12] to placing pebbles on multiple targets. Good resources on this topic include [3, 14, 16, 18], and we list more specific references in Subsections 1.3 and 1.4, below. In this paper, we provide a formula for the target pebbling number of a tree, along with an algorithm for calculating it that runs in polynomial time.
We begin with necessary graph theory and pebbling definitions in Subsections 1.1 and 1.2. In Subsections 1.3 and 1.4 we discuss the relevant history and motivation for our results. Subsection 1.5 contains our two main results regarding the formula and algorithm mentioned above, namely, Theorems 11 and 12. (It takes us time to develop the concepts and notation necessary to be able to state the formula in Corollary 21.) Then we prove these results in Section 2 and conclude in Section 3 with additional comments and questions.
1.1 Graph Definitions
In this paper a graph $G(V,E)$ is simple and connected. We write $V$ in place of $V(G)$ unless different graphs need to be distinguished. The degree ${\sf deg}_{G}(v)$ of a vertex $v$ in a graph $G$, is the number of edges incident to $v$. The indegree ${\sf deg}_{G}^{-}(v)$ of a vertex $v$ in a digraph $G$ equals the number of directed edges $uv$ in $G$, and the outdegree ${\sf deg}_{G}^{+}(v)$ of a vertex $v$ in a digraph $G$ equals the number of directed edges $vu$ in $G$. In a connected graph $G$ the distance from $u$ to $v$, denoted ${\sf dist}_{G}(u,v)$ is the length of the shortest $u,v$-path in $G$. We ignore the subscripts in the above notations when the graph $G$ is understood. The diameter of a graph $G$ is the maximum distance between any pair of vertices $u$ and $v$, ${\sf diam}(G)=\max_{u,v\in V}{\sf dist}(u,v)$.
A tree is a connected acyclic graph. A leaf is a vertex of degree 1; every other vertex is interior. Let the set of leaves of a tree $T$ be denoted by $L(T)$. A graph is chordal if it has no induced cycle of length 4 or more. A simplicial vertex in a graph is a vertex whose neighbors induce a clique. We will denote the set of simplicial vertices by $S(G)$; thus $S(T)=L(T)$ when $T$ is a tree. Chordal graphs are characterized recursively by either being complete or having a simplicial vertex whose removal leaves a chordal graph.
A path partition of a tree $T=(V,E)$ is a set of paths ${\cal P}=\{P_{1},\ldots,P_{\ell}\}$, such that the paths are edge-disjoint and the union of the paths covers all edges of the tree $V(T)=\cup_{i=1}^{\ell}V(P_{i})$. For a multiset $a=\{a_{1},\ldots,a_{m}\}$ of nonnegative integers, we will label its indices so that $a_{1}\geq a_{2}\geq\cdots\geq a_{m}$. For two different multisets $a=\{a_{1}\geq a_{2}\geq\cdots\geq a_{m}\}$ and $b=\{b_{1}\geq b_{2}\geq\cdots\geq b_{m}\}$, we write $a\succ b$ if there is some $j$ such that $a_{i}=b_{i}$ for all $i<j$ and $a_{j}>b_{j}$. When $a\succ b$ we say that $a$ majorizes $b$. A maximum path partition of a tree is a path partition whose multiset of path lengths is not majorized by that of any other path partition. The following algorithm provides maximum path partition ${\cal P}=\{P_{1},\ldots,P_{\ell}\}$ of a rooted tree $(T,r)$ [6]. Set $T_{0}=\{r\}$, define $\ell$ to be the number of leaves in $T$ that are different from $r$, and repeat the following two steps for each $1\leq i\leq\ell$: (a) define $P_{i}$ to be a maximum length path in $T$ with one endpoint in $T_{i-1}$, and (b) define $T_{i}=T_{i-1}\cup P_{i}$.
1.2 Pebbling Definitions
A pebbling function $F$ is any function $F:V\rightarrow{\mathbb{N}}$. We define its support to be ${\sf supp}(F)=\{v\in V\mid F(v)>0\}$, with $s(F)=|{\sf supp}(F)|$. For a pebbling function we write ${\dot{F}}=\{v^{F(v)}\}_{v\in V}$ for its multiset description, and if ${\sf supp}(F)=1$ then we ignore the set braces; furthermore, if $F=\{v^{t}\}$ we write $v^{t}$ instead of $\{v^{t}\}$, and if $t=1$ we write $v$ instead of $v^{1}$. The size of $F$ equals $|F|=\sum_{v\in V}F(v)$. We use the shorthand notation $\min F=\min_{v}F(v)$. A pebbling function $F$ is called positive if $\min F>0$, and stacked if there is a unique $v\in V$ with $F(v)>0$. For any two functions $F:V\rightarrow{\mathbb{N}}$ and $F^{\prime}:V\rightarrow{\mathbb{N}}$ the sum ${\dot{F}}+{\dot{F}}^{\prime}=\{v^{F(v)+F^{\prime}(v)}\}_{v\in V}$. In particular, we can view a vertex $v$ as the pebbling function $\{v\}$, so that the sum $F+v$ is obtained from the function $F$ by increasing the value $F(v)$ by 1. A configuration $C$ is a pebbling function whose value $C(v)$ represents the number of pebbles at vertex $v$. A target $D$ is a pebbling function whose value $D(v)$ represents the demand of pebbles at vertex $v$. The notion of a pebbling function was first introduced in [3]. One point of doing so is to note that both configurations and targets are pebbling functions; albeit playing different roles (i.e., supply versus demand). In this way, a set of unified adjectives and notations can be used for both. Technically, however, it allows one to consider the sum of a configuration and a target, which plays a critical role in the proof of Theorem 13.
For adjacent vertices $u$ and $v$, the pebbling step $u\mapsto v$ consists of removing two pebbles from $u$ and placing one pebble on $v$. For a configuration $C$ and target $D$, we say that $C$ is $D$-solvable (or that there is a $(C,D)$-solution, or that $G$ has a $(C,D)$-solution) if some sequence of pebbling steps places at least $D(v)$ pebbles on each vertex $v$, otherwise $C$ is $D$-unsolvable. For a solution $\sigma$} that solves $\{v_{1},\ldots,v_{k}\}\subseteq{\dot{D}}$, $C[{\sigma}]$ denotes the sub-configuration of pebbles used in $\sigma$. A $(C,D)$-solution ${\sigma}={\sigma}_{1}\ldots{\sigma}_{m}$ is minimal if, for every $i$, ${\sigma}-{\sigma}_{i}$ is not a $(C,D)$-solution. For a $D$-unsolvable configuration $C$, we say that $C$ is $D$-maximal if adding a single additional pebble to any vertex of $G$ yields a $D$-solvable configuration.
A configuration $C$ is called $D$-extremal if it is $D$-maximal of maximum size. The $D$-pebbling number, $\pi(G,D)$, of a graph $G$ is defined to be the minimum number $m$ such that $G$ is $(C,D)$-solvable whenever $|C|\geq m$; that is, it is one more than the size of a $D$-extremal configuration. The $t$-fold pebbling number, $\pi_{t}(G)$, equals $\max_{v\in G}\pi(G,v^{t})$; when $t=1$, we simply write $\pi(G)$. Also, $C$ is solvable if it is $r$-solvable for all $r$.
For a $(C,D)$-solution $\sigma$} and vertex $v\in{\dot{D}}$, we write ${\sigma}_{v}$ for the set of pebbling steps used to solve $v$; thus $\sigma$} can be partitioned into ${\sigma}_{v}$ over all $v\in{\dot{D}}$. A vertex $v$ is big in a configuration $C$ if $C(v)\geq 2$. $B(C)$ denotes the set of big vertices of a configuration $C$. A potential vertex of a configuration $C$ with respect to a target $D$ is any vertex $v\in B(C)\cup({\sf supp}(C)\cap{\dot{D}})$.
A pebbling step $u_{i}\mapsto v_{i}$ is $r$-greedy if ${\sf dist}(v_{i},r)<{\sf dist}(u_{i},r)$, and a sequence of pebbling steps is $r$-greedy if each of its steps is $r$-greedy. A $(C,D)$-solution $\sigma$} is greedy if, for each $v\in{\dot{D}}$, ${\sigma}_{v}$ is $v$-greedy. A configuration $C$ is $D$-greedy if it has a greedy $D$-solution. A graph $G$ is $D$-greedy if for every configuration $C$ with $|C|\geq{\pi}(G,D)$, there is a greedy $(C,D)$-solution. For a $(C,D)$-solution ${\sigma}=\{u_{i}\mapsto v_{i}\}_{i\in I}$, we define its solution subgraph $G_{\sigma}$ to be the digraph with vertices $\cup_{i}\{u_{i},v_{i}\}_{i\in I}$ and arcs $\cup_{i}\{u_{i}v_{i}\}_{i\in I}$. The cost of $\sigma$} equals ${\sf cost}({\sigma})=|I|+|D|$; i.e. the number of pebbles used by $\sigma$}.
1.3 History
The basic lower and upper bounds for every graph are as follows.
Fact 1 ([5, 6]).
For every graph $G$ we have $\max\{n,$ $2^{{\sf diam}(G)}\}\leq\pi(G)\leq(n(G)-{\sf diam}(G))(2^{{\sf diam}(G)}-1)+1$.
The lower bound $\pi(G)\geq n$ is sharp for cubes [6], 2-connected graphs with a dominating vertex, 3-connected graphs of diameter two [13], the Petersen graph, and many others, including, probabilistically, almost all graphs [10] (these are known as Class 0 graphs [13]). The lower bound $\pi(G)\geq 2^{{\sf diam}(G)}$ is sharp for Cartesian products of paths [6], for example. The upper bound in Fact 1 is sharp for paths and complete graphs.
It has been shown that deciding if $C$ is $r$-solvable is NP-complete for general graphs ([15, 18]), as well as for some specific graph classes (e.g., planar or diameter two [8] graphs)), while for others is in P (e.g., planar and diameter two graphs [17], trees [4]). Deciding if ${\pi}(G)\leq K$ takes even longer, in general (in ${\pi}_{2}^{\sf P}$ — see [18]), although formulas for ${\pi}(G)$ that can be calculated in polynomial time are known for many classes of graphs (e.g., Cartesian products of paths [6], trees [4, 6], diameter two graphs [13, 20], sufficiently dense graphs [9], split graphs [2], powers of paths [3], and others — see [14]). In Theorem 12 we provide a polynomial-time algorithm for computing ${\pi}(T,D)$. Some of the motivation for this work lies in the following conjecture. The pyramid is the 6-cycle with a triangle added to one of its maximum independent sets. We say that a graph is pyramid-free if it no induced subgraph is isomorphic to a pyramid.
Conjecture 2 ([3]).
If $G$ is a pyramid-free chordal graph then ${\pi}(G)$ can be calculated in polynomial time.
Suppose $H$ is a spanning subgraph (e.g. a spanning tree) of $G$. Because any $(C,D)$-solution in $H$ is a $(C,D)$-solution in $G$, we have ${\pi}(G,D)\leq{\pi}(H,D)$. Thus the calculation of pebbling numbers of trees is important. Chung [6] used a maximum $r$-path partition of $T$ to calculate the case in which $D=\{r\}$. In this paper we solve the general case (see Corollary 21, below). In both cases the calculation involves a polynomial algorithm that decomposes the structure of an $n$-vertex tree in relation to the target. We state in Theorem 12 that the algorithm runs in $O({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}s(D)\cdot n})$ time.
For the case $D=\{r\}$, Chung used a maximum path partition ${\cal P}=\{P_{1},\ldots,P_{\ell}\}$ of a given rooted tree $(T,r)$. We note that Bunde, et al. [4] proved that a path partition of any tree can be constructed in linear time. Let $l_{i}$ denote the length of the path $P_{i}$, $w_{i}$ be the leaf vertex of $P_{i}$ that is not in $\cup\_{j<i}P_{j}$, and define the configuration ${\hat{C}}_{r,t}$ on $T$ by ${\hat{C}}_{r,t}(w_{1})=t2^{l_{1}}-1$, ${\hat{C}}_{r,t}(w_{i})=2^{l_{i}}-1$ for each $2\leq i\leq\ell$, and ${\hat{C}}_{r,t}(x)=0$ otherwise. We will refer to ${\hat{C}}_{r,t}$ as a Chung configuration. We remark that a Chung configuration is not $t$-fold $r$-solvable — indeed, no pebble on $w_{i}$ can reach the endpoint of $P_{i}$ opposite from $w_{i}$; i.e., no pebble from $P_{i}$ can reach $P_{i-1}$. We note that ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\hat{C}}_{r,t}(w_{1})}=\max{\hat{C}}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}r,t}}$ and observe that the unique $t$-fold $r$-solution of ${\hat{C}}_{r,t}+w_{1}$ uses only the pebbles from $w_{1}$. This property will be a key mechanism in our proofs below. The simplest example of the following theorem is that ${\pi}(P_{n})=2^{n-1}$.
Theorem 3 (Chung [6]).
If $(T,r)$ is a rooted tree then ${\pi}_{t}(T,r)=|{\hat{C}}_{r,t}|+1$.
Thus, every rooted tree $(T,r)$ has a $t$-fold $r$-extremal configuration $C$ with ${\sf supp}(C){\ \subseteq\ }L(T)$. In fact, more can be said. The simple, yet powerful lemma below was first observed by Moews.
Lemma 4 (No-Cycle Lemma [19]).
If $\sigma$} is a minimal $(C,r)$-solution in $G$ then $G_{\sigma}$ is acyclic.
Note that this implies that minimal solutions in trees are greedy. Furthermore, this implies that if $C$ is $t$-fold $r$-unsolvable on a tree $T$, and $C(v)>0$ for some interior vertex $v$, then so is the configuration $C^{\prime}$ defined by $C^{\prime}(v)=C(v)-1$, $C^{\prime}(u)=C(u)+2$ for some neighbor $u$ of $v$ that is farther from $r$, and $C^{\prime}(w)=C(w)$ for all other $w$. In particular, since $|C^{\prime}|>|C|$, this proves the following fact.
Fact 5.
For every rooted tree $(T,r)$, every $t$-fold $r$-extremal configuration $C$ on $T$ has ${\sf supp}(C){\ \subseteq\ }L(T)$.
Later, Alcón and Hurlbert generalized this result to chordal graphs.
Theorem 6 (Alcón, Hurlbert [3]).
If $G$ is chordal then, for all $r\in V$ and $t\geq 1$, there is a $t$-fold $r$-extremal configuration $C$ with ${\sf supp}(C){\ \subseteq\ }S(G)$.
This led to the following conjectured generalization.
Conjecture 7.
If $G$ is chordal and $C$ is $D$-extremal then there is a $D$-unsolvable configuration $C^{*}$ with $|C^{*}|\geq|C|$ and ${\sf supp}(C^{*}){\ \subseteq\ }S(G)$.
This conjecture had been shown for trees, in Theorem 8 of [7], in the case that $D(v)>0$ for all $v\in V(G)$. We prove it for trees and all $D$ in Theorem 11.
1.4 Further Motivation
After Chung generalized the target $r$ to $r^{t}$, the next general target to be considered was $D=V$. In [7], the pebbling number ${\pi}(G,V)$ was called the cover pebbling number. For example, they proved that ${\pi}(P_{n},V)=2^{n}-1$. More generally, they defined the functions ${\alpha}(v,D)=\sum_{u\in V}D(u)2^{{\sf dist}(u,v)}$ for $v\in V$ and ${\alpha}(G,D)=\max_{v\in V}{\alpha}(v,D)$, and proved that, for positive $D$, ${\pi}(T,D)={\alpha}(T,D)$ for every tree $T$. The authors also conjectured that the same formula holds for all graphs, which they proved for complete graphs — the formula simplifies in this case to ${\pi}(K_{n},D)=2|D|-\min D$ for positive $D$. The conjecture was eventually proven, independently, by Vuong and Wyckoff [22] and Sjostrand [21].
Theorem 8 ([21, 22]).
For positive $D$ on any graph $G$, ${\pi}(G,V)={\alpha}(G,V)$.
The main technique in proving this fact is showing that any $D$-extremal configuration is stacked. In general, for non-positive $D$ this is not always true, as we see from Chung configurations on trees. However, for trees with non-positive targets, we are able to generalize the stacking notion to that of a superstack, which will have one “dominant” stack, in the manner of a Chung configuration.
A new tool in attacking the pebbling number of graphs is the Weak Target Conjecture [12].
Conjecture 9 (Weak Target Conjecture).
Every graph $G$ satisfies $\pi(G,D)\leq\pi_{|D|}(G)$ for every target $D$.
Herscovici, et al. [12] showed that the conjecture is true for trees, cycles, complete graphs, and cubes. Hurlbert and Seddiq [16] verified it for the families of 2-paths and Kneser graphs $K(m,2)$. Later, the authors of [3] proposed the more powerful Strong Target Conjecture.
Conjecture 10 (Strong Target Conjecture).
Every graph $G$ satisfies $\pi(G,D)\leq\pi_{|D|}(G)-s(D)+1$ for every target $D$.
They proved that trees and powers of paths [^1] The $k^{\rm th}$ power $G^{(k)}$ of a graph $G$ adds to $G$ all edges $xy$ with ${\sf dist}_{G}(x,y)\leq k$. $P_{n}^{(k)}$ satisfies this stronger conjecture for all $n$ and $k$. In fact, the authors of [3] used these facts to derive the formula for ${\pi}(P_{n}^{(k)})$.
[…]
1.5 Our Results
In this section we present our primary results, including a verification of Conjecture 7 for trees and an algorithm to compute the pebbling numbers ${\pi}(T,D)$ and ${\pi}_{t}(T)$ that runs in polynomial time.
Theorem 11.
If $T$ is a tree and $C$ is $D$-extremal then ${\sf supp}(C){\ \subseteq\ }L(T)$.
We prove Theorem 11 in Subsection 2.2. From this key result, we develop an additional structure that can be imposed on $D$-extremal configurations — see Lemma 20. It is Lemma 20 that enables us to write a formula for ${\pi}(T,D)$ in Corollary 21.
Theorem 12.
For every tree $T$ and target $D$, the values of ${\pi}(T,D)$ and ${\pi}_{t}(T)$ can be computed in $O({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}s(D)\cdot n})$ and $O(n^{2})$ time, respectively.
The proof of Theorem 12 can be found at the end of Section 2, and also rests on the structure imposed by Lemma 20.
2 Proofs
2.1 Paths
Let $F=C+D$ be a pebbling function on a graph. Define ${\cal A}=\langle a_{1},\ldots,a_{m}\rangle$ to be the pebbling arrangement of $F$, with $m=n+|C|+|D|$, where ${\cal A}=\langle v_{1},v_{1,1},\ldots,v_{1,F(v_{1})},v_{2},v_{2,1},\ldots,v_{2,F(v_{2})},\ldots,v_{n},v_{n,1},\ldots,v_{n,F(v_{n})}\rangle$. For example, in Figure 1 we have $G=P_{7}$ with configuration ${\dot{C}}=\{v_{3}^{3},v_{4}^{21},v_{6}^{5}\}$, target ${\dot{D}}=\{v_{1}^{2},v_{2},v_{5},v_{7}^{3}\}$, and pebbling arrangement
${\cal A}=\langle v_{1},v_{1,1},v_{1,2},v_{2},v_{2,1},v_{3},v_{3,1},v_{3,2},v_{3,3},v_{4},v_{4,1},\ldots,v_{4,21},v_{5},v_{5,1},v_{6},v_{6,1},\ldots,v_{6,5},v_{7},v_{7,1},v_{7,2},v_{7,3}\rangle,$
with $m=7+29+7=43$. We leave it to the reader to check that $C$ is a maximal $D$-unsolvable configuration.
Next, let $\cal C$} be the set of maximal $D$-unsolvable configurations $C$ having ${\sf supp}(C)\subseteq\{v_{1},v_{h},v_{n}\}$.
[…]
Figure 1: A path $P$ with both endpoints in $D$, showing the sequence of configurations $C$, $C^{\prime}$, …, and $C_{n}$, (in green), and targets $D$, $D^{\prime}$, …, and $D_{R}^{+}$ (in red) that are used in the proof of Theorem 13.
Theorem 13.
If $P$ is a path and $C$ is $D$-unsolvable then there is a $D$-unsolvable configuration $C^{*}$ with $|C^{*}|\geq|C|$ and ${\sf supp}(C^{*}){\ \subseteq\ }L(P)$.
Proof.
Let $G=P_{n}$, with $|D|=t$. We use induction on $n$ and $t$.
When $n\leq 2$ the statement is trivial, and when $t=1$ the statement is proved by Theorem 6 (since $s(D)=1$). Now we may assume that $n>2$ and $t>1$.
Suppose that some endpoint (without loss of generality, $v_{1}$) is not in $D$. Suppose that the first target is on $v_{k}$, for some $k$, where $1<k<n$. If $C_{[1,k-1]}$ solves $v_{k}$, we use the necessary pebbles to solve $v_{k}$. Let $\sigma$ be a minimum $(C_{[1,k-1]},v_{k})$-solution. $C[{\sigma}]$ is the sub-configuration of pebbles used in $\sigma$. Define $C^{\prime}=C-C[{\sigma}]$ and $D^{\prime}=D-v_{k}$. Then $C^{\prime}$ is not $D^{\prime}$-solvable, and so by induction on $t$ (since $|D^{\prime}|<t$) we know that there exists a $D^{\prime}$-unsolvable configuration $C^{\prime\prime}$, such that $|C^{\prime\prime}|=|C^{\prime}|$ and every interior vertex is empty. Finally, define ${\dot{C}}^{*}={\dot{C}}^{\prime\prime}+v_{1}^{|C[{\sigma}]|}$. Then $C^{*}$ is $D$-unsolvable with every interior vertex empty. If $C_{[1,k-1]}$ doesn’t solve $v_{k}$, we then remove $C_{[1,k-1]}$. Since $C_{[1,k-1]}$ doesn’t solve $v_{k}$ we define $C^{\prime}=C_{[k,n]}$ and set $D^{\prime}=D$. We remove vertices $v_{1}$ through $v_{k-1}$ and obtain a shorter path $P^{\prime}=P_{[k,n]}$; then $C^{\prime}$ is $D^{\prime}$-unsolvable on $P^{\prime}$. We can apply induction on the number of vertices of $P^{\prime}$ to obtain the configuration $C^{\prime\prime}$, which is $D^{\prime}$-unsolvable of size $|C^{\prime}|$ and has ${\sf supp}(C^{\prime\prime})\subseteq\{v_{k},v_{n}\}$. Let ${\dot{C}}^{*}=v_{1}^{|C_{[1,k]}|+C^{\prime\prime}(v_{k})}+({\dot{C}}^{\prime\prime}-v_{k}^{C^{\prime\prime}(v_{k})})$. Since $C^{\prime\prime}$ is $D$-unsolvable, $C^{*}$ is also $D$-unsolvable.
If both endpoints are in $D$ we define ${\cal A}=\langle a_{1},\ldots,a_{m}\rangle$ be the pebbling arrangement (as shown before the statement of Theorem 13 and in Figure 1) of $C+D$, with $m=n+|C|+|D|$. Let $\cal C$} be the set of maximal $D$-unsolvable configurations $C$ having ${\sf supp}(C)\subseteq\{v_{1},v_{h},v_{n}\}$.
Let $j$ be the minimum such that $C_{\langle 1,j\rangle}$ solves $D_{\langle 1,j\rangle}$, and let $\sigma$} denote this solution. Define $h=h(j)$ to be such that each pebble $a_{j,i}$ is on vertex $v_{h}$. (In the example of Figure 1 we have $h=4$, $|C[{\sigma}]|=3+14=17$, and $j=24$, since the $17^{\rm th}$ pebble in $C[{\sigma}]$ is $v_{4,14}$, which is the $24^{\rm th}$ element in $\cal A$}.) Note that $a_{j}\in C[{\sigma}]$, which implies that $D(h)=0$. Hence $1<h<n$. Moreover, the No-Cycle Lemma 4 implies that every step in $\sigma$} moves from right to left (i.e. from some $v_{k}$ to $v_{k-1}$). Therefore, the configuration $v_{h}^{|C[{\sigma}]|}$ either solves $D_{\langle 1,j\rangle}$ exactly (by being equal to $C[{\sigma}]$) or is $D_{\langle 1,j\rangle}$-unsolvable. Define $C^{\prime}=C-C[{\sigma}]$, $D^{\prime}=D-D_{\langle 1,j\rangle}=D_{\langle j+1,m\rangle}=D_{[h,n]}$, and let ${\cal A}^{\prime}=\langle a^{\prime}_{1},\ldots,a^{\prime}_{m^{\prime}}\rangle$ be the pebbling arrangement of $C^{\prime}+D^{\prime}$, where $m^{\prime}=(n-h+1)+|C^{\prime}|+|D^{\prime}|$. (In Figure 1 we have ${\dot{C}}^{\prime}=\{v_{6}^{7}\}$, ${\dot{D}}^{\prime}=\{v_{5},v_{7}^{3}\}$ and
${\cal A}^{\prime}=\langle v_{1},v_{2},v_{3},v_{4},v_{4,1},v_{4,2},\ldots,v_{4,7},v_{5},v_{5,1},v_{6},v_{6,1},\ldots,v_{6,5},v_{7},v_{7,1},v_{7,2},v_{7,3}\rangle$
with $m^{\prime}=7+12+4=23$.) Then $\cal A$$^{\prime}$ is $D^{\prime}$-extremal.
We now have that $|D^{\prime}|<|D|$, so we can use induction on $t$ to know that there is a $D^{\prime}$-unsolvable configuration $C^{\prime\prime}$ on $P_{[h,n]}$ with $|C^{\prime\prime}|=|C^{\prime}|$ and every interior vertex is empty; i.e. ${\sf supp}(C^{\prime\prime})\subseteq\{v_{h},v_{n}\}$. We move $C[{\sigma}]$ to $v_{h}$; i.e. define ${\dot{C}}^{*}={\dot{C}}^{\prime\prime}+v_{h}^{|C[{\sigma}]|}$. (In Figure 1, we have $|C^{\prime\prime}|=|C^{\prime}|=12$, with ${\dot{C}}^{\prime}=\{v_{4}^{12}\}$, ${\dot{D}}^{\prime\prime}={\dot{D}}^{\prime}=\{v_{6},v_{7}^{3}\}$, ${\dot{C}}^{*}={\dot{C}}^{\prime\prime}+v_{h}^{|C[{\sigma}]|}$, and $|C^{*}|=12+17=29$.) We argue by contradiction that $C^{*}$ is $D$-unsolvable. Indeed, suppose that $C^{*}$ is $D$-solvable. Then $D_{\langle 1,j\rangle}$ must use all of $v_{h}^{|C[{\sigma}]|}$, and possibly some pebbles from $C^{\prime\prime}$. We are then left with a sub-configuration of $C^{\prime\prime}$ to solve $D^{\prime\prime}=D_{\langle j+1,m\rangle}$, which is impossible since $C^{\prime\prime}$ is $D^{\prime\prime}$-unsolvable. Hence $C^{*}$ is $D$-unsolvable.
We now let $C_{1}$ and $C_{n}$ be maximal $D$-unsolvable configurations stacked on $v_{1}$ and $v_{n}$, respectively, and prove by convexity that either $|C_{1}|\geq|C^{*}|$ or $|C_{n}|\geq|C^{*}|$. (See Figure 1: we need $2+1\cdot 2+1\cdot 2^{4}+3\cdot 2^{6}=212$ pebbles stacked on $v_{1}$ to solve $D$, so $|C_{1}|=211$; similarly, $|C_{7}|=166$. Observe that $29=|C^{*}|<(|C_{1}|+|C_{7}|)/2$.)
Recall that $\cal C$} is the set of maximal $D$-unsolvable configurations $C$ having ${\sf supp}(C)\subseteq\{v_{1},v_{h},v_{n}\}$, which we observe is nonempty because of the existence of $C^{*}$. We use the convexity of binary exponentiation to prove that the largest configuration in $\cal C$} has no pebbles on $v_{h}$. Let ${\dot{C}}^{+}={\dot{C}}^{*}+v_{1}$; then $C^{*}$ solves $D$ via the solution we call $\sigma$$^{+}$. We denote by $D_{L}$ the multiset of all target vertices $v_{i}$ with $i<h$ that are solved by $\sigma$$^{+}$ using pebbles from $v_{h}$; $D_{R}^{-}$ the multiset of all target vertices $v_{j}$ with $j>h$ that are solved by $\sigma$$^{+}$ using pebbles from $v_{h}$; and $D_{R}^{+}$ the multiset of all target vertices $v_{k}$ with $k>h$ that are solved by $\sigma$$^{+}$ using pebbles from $v_{n}$. (For example, in Figure 1 we have that ${\dot{D}}_{L}=\{v_{1}^{2},v_{2}^{1}\}$, ${\dot{D}}_{R}^{-}=v_{5}$, and ${\dot{D}}_{R}^{+}=v_{7}$.) Then
$|C^{*}|=|C^{+}|-1$ $=|C^{+}(v_{h})|+|C^{+}(v_{n})|-1$
$=\left(\sum_{v_{i}\in D_{L}}2^{h-i}+\sum_{v_{j}\in D_{R}^{-}}2^{j-h}+\sum_{v_{k}\in D_{R}^{+}}2^{n-k}\right)-1.$
As noted above, we know that $1<h<n$. Suppose that some $v_{j}\in D_{R}^{-}$ has jh<njj-h<n-j; i.e. $j<(n+h)/2$. Then define $C^{*^{\prime}}=C^{*}-v_{h}^{2^{j-h}}+v_{n}^{2^{n-j}}$. We see that $C^{*^{\prime}}\in{\cal C}$ and $|C^{*^{\prime}}|>|C^{*}|$, and so we may assume that no such $j$ exists; that is, every $v_{j}\in{\dot{D}}_{R}^{-}$ has $j\geq(n+h)/2$. This implies that every $v_{k}\in{\dot{D}}_{R}^{+}$ has $k\geq(n+h)/2>(n+1)/2$; i.e. k1>nkk-1>n-k. Now we have
$2|C^{*}|$ $=\left(2\sum_{v_{i}\in D_{L}}2^{h-i}+2\sum_{v_{j}\in D_{R}^{-}}2^{j-h}+2\sum_{v_{k}\in D_{R}^{+}}2^{n-k}\right)-2$
$\leq\left(\sum_{v_{i}\in D_{L}}2^{n-i}+\sum_{v_{j}\in D_{R}^{-}}2^{j-1}+\sum_{v_{k}\in D_{R}^{+}}(2^{n-k}+2^{n-k})\right)-2$
$<\left(\sum_{v_{i}\in D_{L}}(2^{i-1}+2^{n-i})+\sum_{v_{j}\in D_{R}^{-}}(2^{j-1}+2^{n-j})+\sum_{v_{k}\in D_{R}^{+}}(2^{k-1}+2^{n-k})\right)-2$
$=|C_{1}|+|C_{n}|.$
Hence $|C^{*}|<(|C_{1}|+|C_{n}|)/2$, and so either $|C_{1}|>|C^{*}|$ or $|C_{n}|>|C^{*}|$. This finishes the proof. ∎
Let $D$ be a target on a graph $G$ with ${\dot{D}}=\{v_{i_{1}},\ldots,v_{i_{t}}\}$ so that $|D|=t$. For a vertex $v$ let ${\alpha}_{G}(v,D)$ be the minimum number such that the configuration $v^{{\alpha}_{G}(v,D)}$ solves $D$ on $G$. (Notice that this is the same function $\alpha$} as used in Theorem 8, although $D$ here is not necessarily positive.) For example $v_{1}^{2^{k}}$ solves $v_{k+1}$ on $P_{k+1}$ but $v_{1}^{2^{k}-1}$ does not, and so ${\alpha}_{P_{k+1}}(v_{1},v_{k+1})=2^{k}$. For ease of reading, we shall write $\alpha$} in place of ${\alpha}_{G}$ when the graph $G$ is understood.
Now let $G=P_{n}$ be the path $v_{1}\cdots v_{n}$. For $1\leq j\leq t$ define on $G$ ${\dot{D}}_{j}^{L}=\{v_{i_{1}},\ldots,v_{i_{j}}\}$ and ${\dot{D}}_{j}^{R}=\{v_{i_{j}}\ldots,v_{i_{t}}\}$. Additionally, define the configuration
${\dot{F}}_{j}=\{v_{1}^{{\alpha}_{P_{n}}(v_{1},{\dot{D}}_{j}^{L})-1},v_{n}^{{\alpha}_{P_{n}}(v_{n},{\dot{D}}_{j}^{R})-1}\};$
that is, the number of pebbles on $v_{1}$ is one shy of the number that would solve all of ${\dot{D}}_{j}^{L}$, and the number of pebbles on $v_{n}$ is one shy of the number required to solve all of ${\dot{D}}_{j}^{R}$. Finally, let $f_{j}=|F_{j}|$ for $1\leq j\leq t$, and define $f(D,n)=\max\{f_{1},f_{t}\}+1$. The following fact is evident.
Fact 14.
Given the target ${\dot{D}}=\{v_{i_{1}},\ldots,v_{i_{t}}\}$ on the path $P_{n}$ with vertex $v$, let $d_{j}={\sf dist}(v,v_{i_{j}})$. Then ${\alpha}_{G}(v,D)=\sum_{j=1}^{t}2^{d_{j}}$.
Lemma 15.
Given the target ${\dot{D}}=\{v_{i_{1}},\ldots,v_{i_{t}}\}$ on the path $P_{n}$, let $f_{k}$ be defined as above. Then the sequence $f_{1},\ldots,f_{t}$ is unimodal; in particular, $f_{h}>f_{h-1}$ when $i_{h}\leq(n+1)/2$ and $f_{h}<f_{h-1}$ when $i_{h}>(n+1)/2$. Therefore $\max_{k}f_{k}=\max\{f_{1},f_{t}\}$.
Proof.
From Fact 14 we obtain the following formula:
$f_{k}=\sum_{j=1}^{h}\left(2^{{\sf dist}(v_{1},v_{i_{j}})}\right)-1+\sum_{j=h}^{t}\left(2^{{\sf dist}(v_{n},v_{i_{j}})}\right)-1.$
From this it follows that $f_{h}-f_{h-1}=2^{{\sf dist}(v_{1},v_{i_{h}})}-2^{{\sf dist}(v_{n},v_{i_{h}})}$, which is nonnegative if and only if ${\sf dist}(v_{1},v_{i_{h}})\geq{\sf dist}(v_{n},v_{i_{h}})$; i.e., $i_{h}-1\geq n-i_{h}$. ∎
Corollary 16.
Suppose that $D$ is a target on the path $P_{n}$. Then $\pi(P_{n},D)=f(D,n)$.
Proof.
Let $D$ be a target with ${\dot{D}}=\{v_{i_{1}},\ldots,v_{i_{t}}\}$, where $i_{1}\leq\dots\leq i_{t}$. For each $1\leq j\leq t$ let $D_{L}(j)=\{v_{i_{1}},...,v_{i_{j}}\}$ and $D_{R}(j)=\{v_{i_{j}},...,v_{i_{t}}\}$. For each $1\leq j\leq t$ let $L_{j}$ be the maximum $D_{L}(j)$-unsolvable stack on $v_{1}$, $R_{j}$ be the maximum $D_{R}(j)$-unsolvable stack on $v_{n}$, and $F_{j}=L_{j}+R_{j}$. By Fact 14 $|L_{j}|={\alpha}(v_{1},D_{L}(j))-1$ and $|R_{j}|={\alpha}(v_{n},D_{R}(j))-1$. For each $1\leq j\leq t$ define $f_{j}=|F_{j}|$. By Lemma 15, among all of the functions $f_{j}$, the maximum value is achieved at one of the endpoints, so $|C|=\max_{j=1}^{t}f_{j}=\max\{f_{1},f_{t}\}$. Suppose that $C$ is a $D$-extremal configuration. By Theorem 13 we may assume that $supp(C)\subseteq\{v_{1},v_{n}\}$. Let $C_{L}=C_{v_{1}}$ and $C_{R}=C_{v_{n}}$. Define $j$ to be the maximum such that $C_{L}$ solves $D_{L}(j-1)$ but not $D_{L}(j)$. Then, by maximality of $C$, $C_{R}$ solves $D_{R}(j+1)$ but not $D_{R}(j)$. Also, by maximality of $C$, $C_{L}=L_{j}$ and $C_{R}=R_{j}$; therefore $C=F_{j}$. Since $C$ is $D$-extremal, $C\in\{F_{1},F_{n}\}$. Hence, $\pi(P_{n},D)=|C|+1=\max\{f_{1},f_{t}\}+1=f(D,n)$. This concludes the proof. ∎
2.2 Trees
Proof of Theorem 11..
Let $T$ be a tree with $n$ vertices with $|D|=t$. If $T$ has $n\leq 3$ or $|L(T)|=2$, then $T$ is a path and the result follows from Theorem 13. So let assume otherwise, that $n\geq 4$ and $|L(T)|\geq 3$. Additionally, if $s(D)=1$, then the result follows from Fact 5, so we assume that $s(D)\geq 2$, which implies that $t\geq 2$.
Let $C$ be a $D$-extremal configuration — we may assume that $C$ has the fewest number of interior pebbles among all such configurations.
Let $P$ be any leaf-split vertex path in $T$, where $x$ is a leaf of $P$ and $y$ be the split vertex (opposite of $x$) in $P$. Define $P^{\prime}=P-y$ and $T^{\prime}=T-P^{\prime}$. Since $C$ is extremal, $C$ is (Dw)(D-w)-solvable for any $w\in D$. Choose $v\in D$ to be the closest to $x$, let $w\in D-v$, and let $\sigma$} be a minimum (C,Dw)(C,D-w)-solution. We now analyze ${\sigma}_{v}$ and define $C^{\prime}=C-C[{\sigma}]$ and $D^{\prime}=D-v$.
Consider the case that ${\sigma}_{v}$ comes from the $x$-side of $v$. If ${\sigma}_{v}$ uses a pebble on an interior vertex $z$, then the configuration that replaces that pebble by a pebble on $x$ is a $D$-extremal configuration with fewer interior pebbles, a contradiction. Hence, $C[{\sigma}]$ is on $x$ already. Since ${\sigma}_{v}$ is minimum, there are no interior pebbles on $P^{\prime}$. Since $C^{\prime}$ is $D^{\prime}$-extremal in $T$ and $|D^{\prime}|<|D|$, induction implies that ${\sf supp}(C^{\prime}){\ \subseteq\ }L(T)$. Hence ${\sf supp}(C){\ \subseteq\ }L(T)$.
Now consider the case that ${\sigma}_{v}$ comes from the $y$-side of $v$. Let $\tau$} be the minimum number of pebbles needed to add to $C^{\prime}$ at $y$ to solve $D$, and define $D^{\prime\prime}$ on $T^{\prime}$ by setting $D^{\prime\prime}(y)=D(y)+{\tau}$ and $D^{\prime\prime}(u)=D(u)$ for all $u\in T^{\prime}-y$. Then $C^{\prime}$ is $D^{\prime\prime}$-extremal. We proceed by induction on the number of leaves since $|L(T^{\prime})|<|L(T)|$ because $y$ is a split vertex in $T$. Hence ${\sf supp}(C^{\prime}){\ \subseteq\ }L(T)$. Therefore ${\sf supp}(C){\ \subseteq\ }L(T)$, which completes the proof. ∎
Let $\cal C$} be the set of all $D$-unsolvable configurations $C$ for which ${\sf supp}(C){\ \subseteq\ }L(T)$. For $C\in{\cal C}$ and $v\in{\sf supp}(C)$, define the stacked configuration $C_{v}$ by $C_{v}(v)=C(v)$ and $C_{v}(x)=0$ otherwise, and also define $C_{v}^{+}$ by $C_{v}^{+}(v)=C_{v}(v)+1$ and $C_{v}^{+}(x)=0$ otherwise. Next we say that $v$ is a superstack if the configuration $C_{v}^{+}$ is $D$-solvable. Now let $C^{*}_{v}$ be any $D$-extremal configuration in $\cal C$} containing $C_{v}$ — which we call a superstack configuration — and define $\cal C$$^{*}$ to be set of $C^{*}_{v}$ for all leaves $v$. Note that when $D=\{r^{t}\}$ we have $C^{*}_{v}={\hat{C}}_{r,t}$.
As an aside, it is worth noting that superstacks do not characterize extremal configurations. Indeed, consider $K_{1,3}$ with leaves $r$, $u$, and $v$, and target $D=r^{2}$. Then $C_{1}=\{u^{7},v\}$ and $C_{2}=\{u^{5},v^{3}\}$ are both $D$-extremal, but only $C_{1}$ is a superstack.
Define a $(C,D)$-solution $\sigma$} to be merging if there are distinct vertices $u$, $v$, and $w$ such that $\sigma$} contains the pebbling steps $u\mapsto w$ and $v\mapsto w$; otherwise, it is merge-free. Note that if $\sigma$} is a merge-free solution of a single target then $T_{\sigma}$ is a path. Such a vertex $w$ is called a merging vertex of $\sigma$}; observe that a merging vertex of $\sigma$} is identified by having ${\sf deg}_{T_{\sigma}}^{-}(v)\geq 2$. For a $(C,D)$-solution $\sigma$}, we define its merge number $\mu({\sigma})=\sum_{v}({\sf deg}_{T_{\sigma}}^{-}(v)-1)$ over all its merging vertices $v$. For a $D$-solvable configuration $C$, we define its merge number $\mu(C)=\min_{\sigma}\mu({\sigma})$ over all $(C,D)$-solutions $\sigma$}.
For a $(C,D)$-solution $\sigma$}, we define the source of $\sigma$} to be ${\mathfrak{S}}({\sigma})=\{v\mid v\in{\sf supp}(C[{\sigma}])\}$. If ${\mathfrak{S}}({\sigma})=\{w\}$ then we say that $w$ is the source of $\sigma$}.
We record the following evident fact without proof.
Fact 17.
Let $a=\{a_{1},\ldots,a_{m}\}$ be a multiset of nonnegative integers, written so that $a_{1}\geq\cdots\geq a_{m}$. Let $1\leq j<j^{\prime}\leq m$ and $c\leq a_{j^{\prime}}$, and define $a^{\prime}=\{a^{\prime}_{1},\ldots,a^{\prime}_{m}\}$ by $a^{\prime}_{j^{\prime}}=a_{j^{\prime}}-c$, $a^{\prime}_{j}=a_{j}+c$, and $a^{\prime}_{i}=a_{i}$ otherwise. Then $a^{\prime}\succ a$. ∎
We will apply this notion to the multiset of values of a configuration $C$: i.e., $\{C(v)\}_{v\in V}$.
Lemma 18 (No-Merging Lemma).
Let $D$ be a target in a tree $T$. Then there is a $D$-extremal configuration $C$, and a leaf with $C(v)=\max C$, such that any $(C+v,D)$-solution $\sigma$} is non-merging.
Observe that paths have these properties by Corollary 16.
55jj11ww77111515zz1111xx114747yy
   55jj11ww77111515zz33xx115555yy
Figure 2: A tree $T$ with target $D$ (in red) and $D$-extremal configurations (in green) $C$ (on the left) and $C^{\prime}$ (on the right), illustrating Lemma 18. Both $w$ and $j$ are merging vertices of the $(C+y,D)$-solution $\sigma$}, with $w$ on the $jy$-path of $T_{\sigma}$. Then the argument in the proof of Lemma 18 converts $C$ to $C^{\prime}$ by moving $2(2^{2})=8$ pebbles from $x$ to $y$.
Proof.
When $s(D)=1$ Lemma 18 follows from Theorem 3, and so we may assume that $s(D)\geq 2$.
Let $C$ be a $D$-extremal configuration on a tree $T$ with $C\succ C^{\prime}$ for all $D$-extremal configurations $C^{\prime}$. By Theorem 11 ${\sf supp}(C)\subseteq L(T)$. Choose any $v$ with $C(v)=\max C$. If $\mu(C+v)=0$ we are done, so we suppose that $\mu(C+v)>0$ and let $\sigma$} be a $(C+v,D)$-solution, where $\mu(C+v)=\mu({\sigma})$. Choose a merging vertex $w$, such that if $u$ is a leaf in $T_{{\sigma}}$, then $w$ is the only merge vertex on the $uw$-path of $T_{{\sigma}}$. Then ${\mathfrak{S}}({\sigma}_{w})=\{x,y\}$. Label $x$ and $y$ so that $C(x)\leq C(y)$. (See Figure 2.)
Suppose that $\sigma$} moves exactly $k$ pebbles from $x$ onto $w$. Then $2\leq k2^{d_{x}}\leq(C+v)(x)<(k+1)2^{d_{x}}$. Define the configuration $C^{\prime}$ by $C^{\prime}(x)=(C+v)(x)-k2^{d_{x}}$, $C^{\prime}(y)=(C+v)(y)+k2^{d_{x}}$, and $C^{\prime}(u)=(C+v)(u)$ for all other $u$. If $C^{\prime}$ is not $D$-solvable, that contradicts $C$ being $D$-extremal, and so $C^{\prime}$ is $D$-solvable, so we assume that $C^{\prime}$ is $D$-solvable. Because $C^{\prime}$ can only contribute an additional $k$ pebbles to $w$, $C^{\prime}-v$ is $D$-unsolvable. Also, $|C^{\prime}-v|=|C|$, and so $C^{\prime}-v$ is $D$-extremal. However, $C^{\prime}-v\succ C$ by Fact 17, which is a contradiction.
Therefore, $\mu(C+v)=0$, which completes the proof. ∎
For vertices $a,b\in T$ denote by $T_{ab}$ the unique $ab$-path in $T$. Define a bush to be an orientation of the edges of a tree with every vertex of indegree at most 1 and a unique vertex $u$ of indegree 0. The vertex of indegree 0 is called the seed.
Thus we may denote the bush with seed $a$ by $B_{a}$.
Corollary 19.
Let $D$ be a target in a tree $T$. Then there is a $D$-extremal configuration $C$, and a leaf $v$ with $C(v)=\max C$, such that the unique $(C+v,D)$-solution $\sigma$} has $|{\mathfrak{S}}({\sigma})|=1$.
Observe that paths have this property by Corollary 16.
Proof.
The No-Merging Lemma 18 implies that, for $v\in L(T)$, with $C(v)=\max C$ the unique $(C+v,D)$-solution $\sigma$} is a pairwise disjoint union of bushes, in which ${\mathfrak{S}}({\sigma})$ is the set of its seeds, necessarily containing $v$. Thus, if $s(D)=1$ then $|{\mathfrak{S}}({\sigma})|=1$. Hence we may assume that $s(D)\geq 2$. We derive the contradiction that if $\sigma$} contains more than one bush then $C$ is not $D$-extremal.
[…]
uu112211xx11ww11331122vv
uu112211xx11ww11331122vv
Figure 3: A tree $T$ with bushes $B_{u}$ (in blue), $B_{x}$ (in green), $B_{w}$ (in brown), and $B_{v}$ (in orange), above, and the tree $T$, redrawn with choice of $\prec$-minimal bush $B_{v}$, below.
Now define $T^{\prime}=B_{u}\cup B_{v}\cup T_{uv}$. Note that $T^{\prime}-E(T_{uv})$ is a disjoint union of bushes $B_{z}$, each with its seed $z\in V(T_{uv})$. Now define the pebbling function $D^{\prime}$ on $T_{uv}$ by $D^{\prime}(z)={\alpha}_{B_{z}}(z,D_{z})$, where $D_{z}=D\cap B_{z}$. (See Figure 4.)
[…]
uu11442244vv
Figure 4: The tree $T^{\prime}=B_{u}\cup B_{v}\cup T_{uv}$ (derived from Figure 3), showing bushes $B_{z_{u}}$ (in blue) and $B_{z_{v}}$ (in orange), with corresponding seeds $z_{u}$ and $z_{v}$, above, and the pebbling function $D^{\prime}$ on $T_{uv}$.
Clearly, a configuration with support in $\{u,v\}$ solves $D_{u}\cup D_{v}$ on $B_{u}\cup B_{v}$ if and only if it solves $D^{\prime}$ on $T_{uv}$. Let $\sigma$$^{\prime}$ be the $(C+v,D^{\prime})$-solution on $T_{uv}$ induced by $\sigma$}. Because $C$ is $D$-extremal on $T$, it is also $D^{\prime}$-extremal on $T_{uv}$. Therefore, by Lemma 15 and Corollary 16 we know that a pebble added to a $D^{\prime}$-extremal configuration on a path consists of a single bush, contradicting that $\sigma$$^{\prime}$ consists of two bushes. ∎
Lemma 20.
If $D$ is a target on a tree $T$ then there is a $D$-extremal configuration $C\in{\cal C}^{*}$.
Proof.
For a target in $D$ we can find a $D$-extremal configuration $C$ and a vertex $v$ with $(C+v,D)$-solution $\sigma$} with $|{\mathfrak{S}}({\sigma})|=1$ by Corollary 19; that is, vertex $v$ is a superstack. By Theorem 11 ${\sf supp}(C)\subseteq L(T)$. Hence, $C\in{\cal C}^{*}$. ∎
Let $f(T,D)=\max_{v\in L(T)}{|C_{v}^{*}|}+1$. The following corollary follows immediately from Lemma 19.
Corollary 21.
Suppose that $D$ is a target on the tree $T$. Then $\pi(T,D)=f(T,D)$. $\Box$
Now we discuss a method for calculating $f(T,D)$. Let $T(D)$ be the convex hull of $D$; that is, it is the smallest subtree of $T$ that contains $D$. Let $V^{\prime}=L(T(D))=\{v_{1},\ldots,v_{j}\}$. For each $i$, $1\leq i\leq j$, let $d^{\prime}_{i}=\sum_{v\in D}D(v)2^{{\sf dist}(v_{i},v)}$. This equals the minimum number of pebbles placed on $v_{i}$ that solves $D$. Let $T_{i}^{-}$ be the union of components of $T-v_{i}$ that are disjoint from $T(D)$, and set $T_{i}=T_{i}^{-}+v_{i}$. For each $i$ let ${\cal P}_{i}$ denote a maximum path partition of the rooted tree $(T_{i},v_{i})$, $C_{i}$ denote its corresponding Chung configuration, and $w_{i}$ be the leaf vertex opposite $v_{i}$ in the longest path of ${\cal P}_{i}$. Then there is some $i$ such that $|C_{w_{i}}^{*}|+1=f(T,D)$. Hence $C_{w_{i}}^{*}=C_{i}-w_{i}^{C_{i}(w_{i})}+w_{i}^{C_{w_{i}}^{+}(w_{i})}$.
Proof of Theorem 12.
Choose any $v\in{\dot{D}}$ and run breadth-first search from $v$ on $T$ to find the shortest paths from $v$ to all other $u\in{\dot{D}}$; the union of those paths equals $T(D)$, whose leaves equal $V^{\prime}=\{v_{1},\ldots,v_{j}\}$, for some $j\leq s(D)$. This can be done in $O(n)$ steps. For each $i\leq j$, run Dijkstra’s algorithm on $T(D)$ to find the distances ${\sf dist}(u,v_{i})$ for all $u\in V^{\prime}$; this calculates $d^{\prime}_{i}$ for each $i$ and takes $O({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}s(D)\cdot n})$ steps. The construction of all path partitions takes $O(n)$ steps (see [4]), and then finding the maximum in the definition of $f(T,D)$ takes $j\leq s(D)$ steps. Hence the entire process takes $O({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}s(D)\cdot n})$ steps. ∎
3 Conclusion
In light of our results for general targets on trees, we propose expanding Conjecture 2 to the following.
Conjecture 22.
If $G$ is a pyramid-free chordal graph, and $D$ is any target, then ${\pi}(G,D)$ can be calculated in polynomial time.
Along these lines, we offer the following proposition that may be of use in pursuit of this conjecture.
Proposition 23.
If $G$ is chordal and $C$ is maximal $D$-unsolvable then $S(G)-{\sf supp}(D){\ \subseteq\ }{\sf supp}(C)$.
Proof.
If some $v\in S(G)-{\sf supp}(D)$ has $C(v)=0$ then define the configuration $C^{\prime}$ by $C^{\prime}(v)=1$ and $C^{\prime}(u)=C(u)$ for all other $u$. Because $C$ is $D$-maximal, $C^{\prime}$ solves $D$; let $\sigma$$^{\prime}$ be a $(C^{\prime},D)$-solution, which we assume to be acyclic by the No-Cycle Lemma. Because $C$ is $D$-unsolvable, $\sigma$$^{\prime}$ contains a step $v\mapsto y$, for some $y$, which requires $\sigma$$^{\prime}$ to also contain a step $x\mapsto v$. Because $\sigma$$^{\prime}$ is acyclic, $x\not=y$. Now define ${\sigma}={\sigma}^{\prime}-(x\mapsto v\mapsto y)+(x\mapsto y)$ and observe that $\sigma$} is a $(C,D)$-solution, which is a contradiction. ∎
Finally, in support of continued pursuit of Conjectures 7, 9, and 10, we offer some open problems that may aid progress towards this goal.
Problem 24.
Prove a No-Merging Lemma for specific families of chordal graphs such as split graphs, interval graphs, $k$-trees, etc.
Problem 25.
Generalize Lemma 20 to specific families of chordal graphs such as split graphs, interval graphs, $k$-trees, etc.